The following is a breakdown of faculty members at IIM-A.
Rank No of males No of females
Professor 20 12
Associate Prof 18 20
Asst Prof 25 30.
What is the probability that a faculty member selected at random is
a) Female.
b) A female professor
c) A female or professor.
d)A professor or asst. Prof
1
Expert's answer
2020-06-14T18:31:31-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & Male & Female & Total \\ \hline
 Professor & 20 & 12 & 32 \\
 \hdashline
 Associate\ Prof & 18 & 20 & 38 \\
 \hdashline
 Assistant \ Prof & 25 & 30 & 55 \\
 \hdashline
 Total & 63 & 62 &125
\end{array}"

a)


"P(F)={62\over 125}=0.496"

b)


"P(F\cap Prof)={12\over 125}=0.096"

c)


"P(F\cup Prof)=P(F)+P(Prof)-P(F\cap Prof)="

"={62\over 125}+{32\over 125}-{12\over 125}={82\over 125}=0.656"

d)


"P(Prof\cup Assist\ Prof)=P(F)+P(Assist \ Prof)="

"={32\over 125}+{55\over 125}={87\over 125}=0.696"